OpenStudy (anonymous):

A bowl of soup at 200° F. is placed in a room of constant temperature of 60° F. The temperature T of the soup t minutes after it is placed in the room is given by T(t) = 60 + 140 e – 0.075 t Find the temperature of the soup 12 minutes after it is placed in the room. (Round to the nearest degree.)

jimthompson5910 (jim_thompson5910):

\[\Large T(t) = 60+140e^{-0.075t}\] \[\Large T(12) = 60+140e^{-0.075*12}\] \[\Large T(12) = 60+140e^{-0.9}\] \[\Large T(12) = 60+140(0.4065696597)\] I'll let you finish.

OpenStudy (anonymous):

Thank you, I got 116.92 degrees F

jimthompson5910 (jim_thompson5910):

then you round that to the nearest degree to get the final answer of 117 degrees F

OpenStudy (anonymous):

Yes, Thank you